# Omni-Video 2: Scaling MLLM-Conditioned Diffusion for Unified Video Generation and Editing

### v1

@@ 1 Introduction
Unified modeling [tong2024metamorph, ma2025janusflow, deng2025bagel, luo2025univid, chen2025blip3, deng2025emerging] has recently emerged as a powerful paradigm for building general-purpose systems capable of understanding and generating multimodal content, including text, images, video, and their combinations. In the realm of unified video modeling, a growing body of work explores joint video generation and editing by coupling multimodal large language models (MLLMs) models with diffusion-based generative backbones [wei2025univideo], while other efforts aim to unify heterogeneous video editing tasks via in-context conditioning [ye2025unic, ju2025editverse]. Despite this progress, building a practical unified video model remains highly nontrivial: Video modeling incurs substantial computational overhead and imposes strict temporal consistency requirements, rendering large-scale training and inference particularly challenging. More importantly, training high-quality video generation models from scratch is prohibitively expensive, typically requiring massive datasets, long training schedules, and specialized infrastructure. As a result, most practical unified video systems must rely on strong pretrained text-to-video diffusion models as their generative foundation, making effective reuse and preservation of such priors a central concern.
@@ 5.2 Video Editing Performance
Table 2: Quantitative comparison results of text-to-video generation methods on the VBench [huang2024vbench] benchmark. Omni-Video 2 achieves strong performance, outperforming leading text-to-video generation methods across a majority of evaluation dimensions. These results show that our framework for enabling complex editing does not compromise the foundational model’s core generative capabilities.
@@
EasyAnimateV5.1 [fu2024easyanimate] 83.42 85.03 77.01 98.00 97.41 99.19 98.02 57.15 69.48 68.61
MiniMax-Video-01 [minimax2024video] 83.41 84.85 77.65 97.51 97.05 99.10 99.22 64.91 63.03 67.17
@@
CogVideoX1.5-5B [yuan2024cogvideox] 82.17 82.78 79.76 96.87 97.35 98.88 98.31 50.93 62.79 65.02
@@
CogVideoX1.5-5B [yuan2024cogvideox] 87.47 69.65 97.20 87.55 80.25 52.91 24.89 25.19 27.30

### v2

@@ 1 Introduction
Unified modeling [tong2024metamorph, ma2025janusflow, deng2025bagel, luo2025univid, chen2025blip3] has recently emerged as a powerful paradigm for building general-purpose systems capable of understanding and generating multimodal content, including text, images, video, and their combinations. In the realm of unified video modeling, a growing body of work explores joint video generation and editing by coupling multimodal large language models (MLLMs) models with diffusion-based generative backbones [wei2025univideo], while other efforts aim to unify heterogeneous video editing tasks via in-context conditioning [ye2025unic, ju2025editverse]. Despite this progress, building a practical unified video model remains highly nontrivial: Video modeling incurs substantial computational overhead and imposes strict temporal consistency requirements, rendering large-scale training and inference particularly challenging. More importantly, training high-quality video generation models from scratch is prohibitively expensive, typically requiring massive datasets, long training schedules, and specialized infrastructure. As a result, most practical unified video systems must rely on strong pretrained text-to-video diffusion models as their generative foundation, making effective reuse and preservation of such priors a central concern.
@@ 5.2 Video Editing Performance
Table 2: Quantitative comparison results of text-to-video generation methods on the VBench [huang2024vbench] benchmark. Omni-Video 2 achieves strong performance, outperforming leading text-to-video generation methods across a majority of evaluation dimensions. These results show that our framework for enabling complex editing does not compromise the foundational model’s core generative capabilities. Quantitative results are quoted from the official Vbench Leaderboard.
@@
EasyAnimateV5.1 [xu2025easyanimate] 83.42 85.03 77.01 98.00 97.41 99.19 98.02 57.15 69.48 68.61
MiniMax-01 [minimax2024video] 83.41 84.85 77.65 97.51 97.05 99.10 99.22 64.91 63.03 67.17
@@
CogVideoX1.5-5B [yang2024cogvideox] 82.17 82.78 79.76 96.87 97.35 98.88 98.31 50.93 62.79 65.02
@@
CogVideoX1.5-5B [yang2024cogvideox] 87.47 69.65 97.20 87.55 80.25 52.91 24.89 25.19 27.30